# Incorporating Procedural Fairness in Flag Submissions on Social Media Platforms

### v1

@@ 1. Introduction
Flagging constitutes the most visible means through which social media users can directly influence platforms’ content moderation decisions. It serves as an ex post moderation measure (Grimmelmann, 2015) that lets users report any post or account they deem inappropriate. Platforms maintain complex sociotechnical systems composed of automated mechanisms and human reviewers (Gorwa et al., 2020; Roberts, 2016; Jhaver et al., 2019b) that continually review flagged posts and, when warranted, impose sanctions, such as removing, down-ranking, or shadow-banning the flagged posts  (Goldman, 2021; Grimmelmann, 2015). Given the large volume of posts submitted to popular platforms (Huang et al., 2024) and the expenses associated with content review (Li et al., 2022; Roberts, 2016), platforms heavily rely on flags to identify and regulate norm violations and maintain the usability of their sites.
@@
Flags are now ubiquitously available as a feature across all social media platforms, but their description, placement, and implementation vary widely. For example, Reddit offers a ‘report’ button below each post, which triggers a one-step selection of the post’s rule violation among 14 categories, including “Hate,” “Sharing personal information,” and “Non-consensual intimate media.” In contrast, each post on X has a ‘Report post’ option embedded in a drop-down menu at the post’s top-right corner; selecting this option prompts a multi-step information-gathering process about what is inappropriate in that post. These design choices constrain and influence how flaggers can articulate their objections about inappropriate posts. Yet, little prior empirical research has examined how differences in flag implementations affect users’ experiences of flag submissions. Since the regular contributions of flag submitters are indispensable for most content moderation systems (Crawford and Gillespie, 2016; Gillespie, 2018a), it is vital that platforms design flags in ways that satisfy flaggers and encourage reporting of inappropriate posts.
@@
We begin to fill this crucial research gap by examining how the design of four commonly deployed flagging components influences end-users. These components include classification schemes, which aid users in specifying their reason for the report through selecting a rule violation category; community guidelines, which direct users and moderators alike in assessing a post’s compliance with platform standards; a text box, which gives users space to detail specifics of their objections; and moderator type, which offers users some insight into who is administering the flag review process. Current implementations of flagging mechanisms usually adopt one or more of these components to enhance the accuracy and perceived validity of reporting. Therefore, we examine how various combinations of these design elements affect how users experience the flag submission process.
@@
Specifically, this paper explores users’ perception of procedural fairness in their interactions with flagging mechanisms. The procedural fairness perspective argues that the legitimacy of a decision-making system derives from public confidence in the fairness of the processes through which decision-makers exercise their authority  (Sunshine and Tyler, 2003; Tyler, 2006; Vaccaro et al., 2020). It has been widely argued that procedural fairness is a pivotal principle that should inform the design of social media governance systems (Schoenebeck et al., 2021; Suzor, 2019; Matthew Katsaros and Tyler, 2024; Hartmann, 2020; Jhaver et al., 2019a). Enhanced perceptions of procedural fairness foster users’ support for moderation decisions  (Valcke et al., 2020; Katsaros et al., 2022a; Jhaver et al., 2019a) and correlate with norm-following behavior (Tyler et al., 2021), thereby strengthening trust between users and the platform  (Jhaver et al., 2019a, c) and bolstering the legitimacy of governance mechanisms  (Fan and Zhang, 2020; Gonçalves et al., 2021). Prior research on enacting procedural fairness in content moderation systems has examined how moderation outcomes are delivered to and experienced by end-users (Schoenebeck et al., 2021; Ma et al., 2023a; Suzor et al., 2019; Jhaver et al., 2019a). Building upon this rich literature, we concentrate on incorporating procedural fairness in flagging mechanisms from the perspective of flag submitters.
@@
Consistency refers to the level of uniformity of the content moderation process, i.e., how content moderation policies are enacted to make moderation decisions regardless of the variable posting contexts (Vaccaro et al., 2020; Katsaros et al., 2022a). Prior research shows that Black, Indigenous, and people of color (BIPOC), LBGTQIA+, disabled, plus-sized individuals, and sex workers experience more frequent and severe moderation consequences than other users  (Haimson et al., 2021; Ma and Kou, 2022; Are, 2023; Lyu et al., 2024) despite being subjected to the same policies. Such inequitable moderation experiences have contributed to user perceptions that moderation has become a form of censorship (Ma and Kou, 2022).
Our study also addresses disproportionate moderation experiences, but we focus on a different stakeholder — users who flag content. We posit that users’ belief in the platform’s ability to enact consistent flag reviews, regardless of the flagger’s identity, is crucial. Thus, prioritizing inter-flagger “consistency” would mitigate disproportionality in the review process and promote a perception of moderation fairness among users.
We examine factors that shape consistency by exploring the following research question:
@@
Transparency enables individuals to see how decisions are made and ensure social accountability (Chan Kim and Mauborgne, 1998; Breton et al., 2007). Enacting transparency in decision-making positively influences users’ perception of the decision-making system, even when the decisions are unfavorable (Suzor et al., 2019; Ma et al., 2023a; Thach et al., 2022; Brockner and Wiesenfeld, 2013). In the context of content moderation, transparency has been conceptualized as the communicative steps that platforms may take to better explain the deeper complexities of moderation processes, policies, and outcomes to the many stakeholders implicated in them (Suzor et al., 2019).
Content moderation scholars have raised concerns about the lack of transparency in moderation systems, citing insufficient details provided to sanctioned users about what rules they have violated and whether their content was removed by an algorithm or a human moderator (Díaz and Hecht-Felella, 2021). Studying the design of personal moderation tools that platforms offer,  Jhaver et al. (2023) highlighted that these tools do not effectively communicate the meanings of their interface elements and offer inadequate information about the inner logic and inputs to algorithms that drive them. Researchers have proposed various ways to enhance transparency in content moderation systems, including providing more details about rules and decision-making processes (Katsaros et al., 2022a; Jhaver et al., 2019a, c), describing how rule violations are detected  (Suzor et al., 2019) and clarifying how sanctions are determined for rule-violating posts (Chandrasekharan et al., 2019).
Given that flagging is a key mechanism that drives content moderation systems, ensuring transparency about it is crucial for maintaining overall system fairness, especially since it can serve as a basis for explaining flagging outcomes. Our study focuses on enacting transparency during the flag submission process, specifically the level of detail available or implied throughout the flagging process about flag reviews. We conceptualize “transparency” as platforms using flagging components to inform flag submitters how flag reviews are processed. We pursue the following research question:
@@
Voice refers to the extent to which users’ input is accommodated in the decision-making process (Katsaros et al., 2022a; Vaccaro et al., 2020). It has been identified as crucial to shaping individuals’ fairness perceptions (Drahos, 2017). When the decision-making process invites users’ input, they are more likely to feel that they have been treated fairly—even when the outcome is unfavorable to them (Thibaut, 1975)—and they perceive greater satisfaction with the process outcome (Van den Bos et al., 1998). In the context of content moderation research,  Ma and Kou (2022) found that YouTube creators associate the fairness of their moderation experiences with having their voice involved in the decision-making process.
Despite its crucial role, limited research examines the perception of voice/agency that users experience when interacting with flagging mechanisms. We conceptualize “voice” in the context of this paper as the users’ perceived degree of involvement in the flagging process, especially regarding how well they can express their objections to an inappropriate post. We thus pursue the following research question:
@@
Transparency Platform provides users with information about moderation procedures and reasoning behind decision-making outcomes (Jhaver et al., 2019a; Ma et al., 2023a; Thach et al., 2022; Suzor et al., 2019; Katsaros et al., 2022a). Flagging mechanism presents relevant information about the flagging procedures and flag reviews.
@@
Voice Platform moderation process adopts measures to integrate users’ input into the decision-making process  (Vaccaro et al., 2020; Katsaros et al., 2022a; Ma and Kou, 2022). Flagging mechanism allows users to thoroughly express their objections to the flagged posts.
@@
To answer the preceding research questions, we conducted a controlled online experiment to evaluate the impact of different flagging implementations on users’ fairness perceptions. We recruited 2,936 participants using Lucid Theorem [^1] https://lucidtheorem.com and randomly assigned each to experience one of 54 flagging scenarios simulated on Qualtrics. These scenarios were constructed by combining different implementations (or absence) of the four flagging components — rule violation classification scheme, guidelines, a text box, and information about the moderator (detailed in Sec. 3, where we motivate our specific research hypotheses).
After reporting a post in such simulated flagging scenarios, participants assessed the procedural fairness of their flagging experience along the dimensions of consistency, transparency, and voice. Next, they answered open-ended questions that solicited their suggestions on how flagging systems could improve procedural fairness.
Following data collection, we conducted a mixed-methods analysis to assess the relationships between the various design choices of flagging components and users’ perceptions of fairness. Additionally, we evaluated how these different components influence flaggers’ cognitive burden of submitting flags and their propensity to flag again.
@@
Our statistical analyses show that displaying community guidelines during the flag submission process raises users’ perceptions of transparency. Additionally, we found that offering a text box where users can input their specific objections about the flagged post enhances their perception of having a voice in that process. Our qualitative analysis of respondents’ open-ended suggestions to enhance flags’ procedural fairness highlights users’ interest in receiving information regarding how the flag review process works, whether the flag reviewers are politically neutral, and how flaggers can offer platforms feedback about their flagging experience. Participants wanted flagging mechanisms to support greater expressivity, e.g., by allowing flaggers to highlight sections of flagged posts that violate platform guidelines or offering alternative channels to submit flags, such as chat, emails, or phone calls. Finally, participants wanted to receive timely notifications of their flag outcomes, and they expected platforms to prevent the abuse of flags by bad actors wishing to take down otherwise appropriate content.
@@
Our findings underscore the necessity for flagging systems to recognize and accommodate users’ nuanced objections about why flagged posts deserve regulation. To achieve this, we recommend that flag implementations (1) integrate a text box for users to input their detailed perspectives to the platform, (2) enhance flags’ vocabulary of complaints (Crawford and Gillespie, 2016) by letting users highlight norm-violating portions of flagged posts and rate the violation severity, and (3) incorporate information and visualization systems that allow flaggers to track the review status of submitted flags. Additionally, platforms should (4) address flaggers’ concerns regarding biased decision-making and false flagging by offering comprehensive information about posting guidelines, reviewers, and the review process. Our empirical insights suggest that platforms could (5) reduce the burden of flagging and flag reviews by engaging in educational efforts to reform the authors of flagged posts, strengthening ex ante moderation measures (Grimmelmann, 2015), and positively reinforcing constructive contributions.
Finally, we demonstrate that users evaluate the fairness of flags not in isolation but by contextualizing flagging within a broader set of platform-offered features that address harm, such as muting and blocking. Thus, platforms must clarify the distinctions between these tools regarding their affordances and required user inputs.
@@ 2.1. Content Moderation
While moderator decisions are made on the platform side, flagging serves as a mechanism for users to participate in the governance process (Chipidza and Yan, 2022). Appeal mechanisms  (Vaccaro et al., 2020, 2021), which let users express their dissatisfaction with content moderation decisions and request that they be reversed, also allow users to communicate with platform operators. However, this mechanism only concerns moderated users and their sanctioned posts, leaving those choosing to flag posts with few or no options to appeal platforms’ decisions. Another bottom-up moderation mechanism that many platforms offer is personal moderation (Jhaver et al., 2023), which lets users limit the visibility of undesirable posts on their feeds. Personal moderation includes actions like muting or blocking an account and customizing the sensitivity threshold for content on one’s feed according to one’s preferences (Jhaver et al., 2023; Jhaver and Zhang, 2023). It exclusively affects one person’s feed without influencing others’ content consumption (Goldman, 2021). For instance, users prioritizing free speech can opt for lenient moderation strategies by increasing the visibility levels of potentially offensive posts in their feeds (Jhaver and Zhang, 2023). By offering the flexibility to adjust their own moderation threshold, personal moderation can balance platforms’ efforts to uphold users’ free speech values while safeguarding them from harm (Chipidza and Yan, 2022).
Prior research examined the design choices involved in building personal moderation tools and recommended improvements in defining interface elements, incorporating cultural context, offering greater granularity, and leveraging example content (Jhaver et al., 2023). We add to this research by evaluating the design space of the flagging mechanism, another crucial bottom-up moderation measure that platforms offer.
@@ 2.2. Flagging as a Content Moderation Tool
Zhang et al. (2023) identified three temporary distinct stages associated with the use of flagging mechanisms — before, during, and after flagging — and highlighted the various user needs and interface affordances available in each stage. Our research focuses on the “during flagging” stage, and examines the design opportunities within that stage to improve users’ flagging experiences.
Platforms would prefer that users flag only content that violates their existing guidelines to optimize the labor involved in content review. As such, platforms seek to concentrate flaggers’ attention on reporting authentic violations of platform guidelines. They do this by designing flagging mechanisms in ways that emphasize their criteria for reviewing flagged content. For instance, flagging interfaces often prompt users to select from predefined categories of rule violations or present platform guidelines that may direct flaggers’ attention to platform norms and serve as an educational tool, making users aware of what platforms consider non-normative behaviors (Chandrasekharan et al., 2018; Matias, 2019b). On some platforms, flagging requires users to articulate their objections to the flagged post in a text box. This encourages users to reflect on the purpose of their flagging and elaborate on their perspectives regarding how the flagged post violates platform norms.
@@
By incorporating these diverse elements, current social media sites demonstrate varied approaches to structuring their flagging mechanisms. For example, Instagram’s flagging system lets users categorize rule violations and specify which rules the reported post has breached (Figure 1). Facebook offers a similar categorization interface and additionally prompts users to review the site’s community guidelines when selecting a rule violation category (Figure 1). YouTube, on the other hand, provides a text box, enabling users to articulate their reasons for flagging in their own words (Figure 1).

### v2

@@ 1. Introduction
Flagging constitutes the most visible means through which social media users can directly influence platforms’ content moderation decisions. Platforms maintain complex sociotechnical systems composed of automated mechanisms and human reviewers (Gorwa et al., 2020; Roberts, 2016; Jhaver et al., 2019b) that continually review flagged posts and, when warranted, impose sanctions, such as removing, down-ranking, or shadow-banning the flagged posts  (Goldman, 2021; Grimmelmann, 2015). Given the expenses associated with content review (Li et al., 2022; Roberts, 2016), platforms rely on flags to identify and regulate norm violations and maintain the usability of their sites.
@@
Flags are now ubiquitously available as a feature across all social media platforms, but their description, placement, and implementation vary widely. For example, Reddit offers a ‘report’ button below each post, which triggers a one-step selection of the post’s rule violation among 14 categories, including “Hate,” “Sharing personal information,” and “Non-consensual intimate media.” In contrast, each post on X has a ‘Report post’ option embedded in a drop-down menu at the post’s top-right corner; selecting this option prompts a multi-step information-gathering process about what is inappropriate in that post. These design choices constrain and influence how flaggers can articulate their objections about inappropriate posts. Yet, little prior empirical research has examined how differences in flag implementations affect users’ experiences of flag submissions. We address this gap by focusing on social media users in their capacity as flag submitters (flaggers). Prior research has shown that the regular contributions of flaggers are indispensable for most content moderation systems (Crawford and Gillespie, 2016; Gillespie, 2018a) — their distributed labor collectively helps moderators address the daunting task of regulating vast and evolving content (Grimmelmann, 2015). Thus, it is vital that platforms design flags in ways that satisfy flaggers and encourage reporting of inappropriate posts.
@@
We begin to fill this crucial research gap by examining how the design of four commonly deployed flagging components influences end-users. These components include classification schemes, which aid users in specifying their reason for the report through selecting a rule violation category; community guidelines, which direct users and moderators alike in assessing a post’s compliance with platform standards; a text box, which gives users space to detail specifics of their objections; and moderator type, which offers users some insight into who is administering the flag review process. Current implementations of flagging mechanisms usually adopt one or more of these components  (Crawford and Gillespie, 2016) to enhance the accuracy and perceived validity of reporting. Therefore, we examine how various combinations of these design elements affect how users experience the flag submission process.
@@
Specifically, this paper explores users’ perception of procedural fairness in their interactions with flagging mechanisms. The procedural fairness perspective argues that the legitimacy of a decision-making system derives from public confidence in the fairness of the processes through which decision-makers exercise their authority  (Sunshine and Tyler, 2003; Tyler, 2006; Vaccaro et al., 2020). It has been widely argued that procedural fairness is a pivotal principle that should inform the design of social media governance systems (Schoenebeck et al., 2021; Suzor, 2019; Matthew Katsaros and Tyler, 2024; Hartmann, 2020; Jhaver et al., 2019a). Prior research on enacting procedural fairness in content moderation systems has examined how moderation outcomes are delivered to and experienced by end-users (Schoenebeck et al., 2021; Ma et al., 2023a; Suzor et al., 2019; Jhaver et al., 2019a). Building upon this rich literature, we concentrate on incorporating procedural fairness in flagging mechanisms from the perspective of flaggers.
@@
Consistency refers to the level of uniformity of the content moderation process, i.e., how content moderation policies are enacted to make moderation decisions regardless of the variable posting contexts (Vaccaro et al., 2020; Katsaros et al., 2022a). Prior research shows that Black, Indigenous, and people of color (BIPOC), LBGTQIA+, disabled, plus-sized individuals, and sex workers experience more frequent and more severe moderation consequences than other users  (Haimson et al., 2021; Ma and Kou, 2022; Are, 2023; Lyu et al., 2024) despite being subjected to the same policies. Such inequitable moderation experiences have contributed to user perceptions that moderation systems discriminate against certain user groups based on their identity characteristics (Thach et al., 2022), further fueling the criticism that moderation has become a form of censorship (Ma and Kou, 2022).
Our study also addresses disproportionate moderation experiences, but we focus on a different stakeholder — users who flag content. We posit that users’ belief in the platform’s ability to enact consistent flag reviews, regardless of the flagger’s identity, is crucial. Thus, prioritizing inter-flagger “consistency” would mitigate disproportionality in the review process and promote a perception of moderation fairness among users. Importantly, platforms should not just enact inter-flagger consistency in their flag review processes but also satisfy end-users that such consistency exists. We examine factors that shape consistency perceptions by exploring the following research question:
@@
Note that while we focus on inter-flagger consistency, many other aspects of flagging consistency warrant attention but are beyond the scope of this study. For example, consistency regarding flag review outcomes for flagged users with different identity characteristics is important, but we chose not to examine it here because we are not concerned with flagged users in this study.
Transparency enables individuals to see how decisions are made and ensures social accountability (Chan Kim and Mauborgne, 1998; Breton et al., 2007). In the context of content moderation, transparency has been conceptualized as the communicative steps that platforms may take to better explain the deeper complexities of moderation processes, policies, and outcomes to the many stakeholders implicated in them (Suzor et al., 2019). Content moderation scholars have raised concerns about the lack of transparency in moderation systems, citing insufficient details provided to sanctioned users about what rules they have violated and whether their content was removed by an algorithm or a human moderator (Díaz and Hecht-Felella, 2021). Researchers have proposed various ways to enhance transparency in content moderation systems, including providing more details about decision-making processes (Katsaros et al., 2022a; Jhaver et al., 2019a, c), describing how rule violations are detected  (Suzor et al., 2019) and clarifying how sanctions are determined for rule-violating posts (Chandrasekharan et al., 2019).
Given that flagging is a key mechanism that drives content moderation systems, ensuring transparency about it is crucial for maintaining overall system fairness, especially since it can serve as a basis for explaining flagging outcomes. Our study focuses on enacting transparency during the flag submission process, specifically the level of detail available or implied throughout the flagging process about flag reviews. We conceptualize “transparency” as platforms using flagging components to inform flaggers how flag reviews are processed. We pursue the following research question:
@@
Voice refers to the extent to which users’ input is accommodated in the decision-making process (Katsaros et al., 2022a; Vaccaro et al., 2020). It has been identified as crucial to shaping individuals’ fairness perceptions (Drahos, 2017). In the context of content moderation research,  Ma and Kou (2022) found that YouTube creators associate the fairness of their moderation experiences with having their voice involved in the decision-making process. Despite its crucial role, limited research examines the perception of voice/agency that users experience when interacting with flagging mechanisms. We conceptualize “voice” in the context of this paper as the users’ perceived degree of involvement in the flagging process, especially regarding how well they can express their objections to an inappropriate post. We thus pursue the following research question:
@@
Transparency Platform provides users with information about moderation process and reasoning behind decision-making outcomes (Jhaver et al., 2019a; Ma et al., 2023a; Thach et al., 2022; Suzor et al., 2019; Katsaros et al., 2022a). Flagging mechanism presents relevant information about the flagging process and flag reviews.
@@
Voice Platform moderation process adopts measures to integrate users’ opinion into the decision-making process  (Vaccaro et al., 2020; Katsaros et al., 2022a; Ma and Kou, 2022). Flagging mechanism allows users to thoroughly express their objections to the flagged posts.
Besides the four flagging components indicated above, we aimed to more deeply understand the changes that end-users seek in the broader design and policy choices associated with flag implementations. Therefore, we ask:
RQ 4: How should flagging processes change to enhance users’ fairness perceptions?
@@
To answer the preceding research questions, we conducted a controlled online experiment to evaluate the impact of different flagging implementations on users’ fairness perceptions. We recruited 2,936 participants using Lucid Theorem [^1] https://lucidtheorem.com and randomly assigned each to experience one of 54 flagging scenarios simulated on Qualtrics. These scenarios were constructed by combining different implementations (or absence) of the four flagging components — rule violation classification scheme, guidelines, a text box, and information about the moderator (detailed in Sec. 3, where we motivate our specific research hypotheses). After reporting a post in such simulated flagging scenarios, participants assessed the procedural fairness of their flagging experience along the dimensions of consistency, transparency, and voice.
@@
Our statistical analyses (RQ 1, 2, 3) show that displaying community guidelines during the flag submission process raises users’ perceptions of transparency. Additionally, we found that offering a text box where users can input their specific objections about the flagged post enhances their perception of having a voice in that process. Our qualitative analysis of respondents’ open-ended suggestions (RQ 4) to enhance flags’ procedural fairness highlights users’ interest in receiving information regarding how the flag review process works and whether the flag reviewers are politically neutral. Participants wanted flagging mechanisms to support greater expressivity, provide timely notifications of flag outcomes, and prevent the abuse of flags by bad actors wishing to take down otherwise appropriate content.
@@
Our findings underscore the need for flagging systems to accommodate users’ nuanced objections about why flagged posts deserve regulation. To achieve this, we recommend that flag implementations (1) integrate a text box for users to input their detailed perspectives to the platform, (2) enhance flags’ vocabulary of complaints (Crawford and Gillespie, 2016) by letting users highlight norm-violating portions of flagged posts and rate the violation severity, and (3) incorporate information and visualization systems that let flaggers track the review status of submitted flags. Additionally, platforms should (4) address flaggers’ concerns regarding biased decision-making and false flagging by offering comprehensive information about posting guidelines, reviewers, and the review process.
@@ 2.1. Content Moderation
While moderation decisions are made on the platform side, flagging serves as a mechanism for users to participate in the governance process (Chipidza and Yan, 2022). Appeal mechanisms  (Vaccaro et al., 2020, 2021), which let users express their dissatisfaction with content moderation decisions and request that they be reversed, also allow users to communicate with platform operators. However, this mechanism only concerns moderated users and their sanctioned posts, leaving those choosing to flag posts with few or no options to appeal platforms’ decisions.
Another bottom-up moderation mechanism that many platforms offer is personal moderation (Jhaver et al., 2023), which lets users limit the visibility of undesirable posts on their feeds. Personal moderation includes actions like muting or blocking an account and customizing the sensitivity threshold for content on one’s feed according to one’s preferences (Jhaver et al., 2023; Jhaver and Zhang, 2023). It exclusively affects one person’s feed without influencing others’ content consumption (Goldman, 2021). Prior research examined the design choices involved in building personal moderation tools and recommended improvements in defining interface elements, incorporating cultural context, offering greater granularity, and leveraging example content (Jhaver et al., 2023). We add to this research by evaluating the design space of the flagging mechanism, another crucial bottom-up moderation measure that platforms offer.
@@ 2.2. Flagging as a Content Moderation Tool
Zhang et al. (2023) identified three temporally distinct stages associated with the use of flagging mechanisms — before, during, and after flagging — and highlighted the various user needs and interface affordances available in each stage. Our research focuses on the “during flagging” stage, and examines the design opportunities within that stage to improve users’ flagging experiences.
Platforms would prefer that users flag only content that violates their existing guidelines to optimize the labor involved in content review. As such, platforms seek to concentrate flaggers’ attention on reporting authentic violations of platform guidelines. They do this by designing flagging mechanisms in ways that emphasize their criteria for reviewing flagged content. For instance, flagging interfaces often prompt users to select from predefined categories of rule violations or present platform guidelines that may direct flaggers’ attention to what platforms consider non-normative behaviors (Chandrasekharan et al., 2018; Matias, 2019). On some platforms, flagging requires users to articulate their objections to the flagged post in a text box. This encourages users to reflect on the purpose of their flagging and elaborate on their perspectives regarding how the flagged post violates platform norms.
@@
By incorporating these diverse elements, current social media sites demonstrate varied approaches to structuring their flagging mechanisms. For example, Instagram’s flagging system lets users categorize rule violations and specify which rules the reported post has breached (Figure 1). Facebook offers a similar categorization interface and additionally prompts users to review the site’s community standards when selecting a rule violation category (Figure 1). YouTube, on the other hand, provides a text box, enabling users to articulate their reasons for flagging in their own words (Figure 1).